OpenStudy (anonymous):

What is the simplified expression for 3 to the power of negative 4 multiplied by 2 to the power of 3 multiplied by 3 to the power of 2 whole over 2 to the power of 4 multiplied by 3 to the power of negative 3?

OpenStudy (anonymous):

Please help me I don't know what the answer is.

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

impossible to read, use the equation tool

OpenStudy (whpalmer4):

\[\frac{3^{-4}*2^3*3^2}{2^4*3^{-3}}\] is that what you are trying to simplify?

OpenStudy (whpalmer4):

I would use the following properties of exponents to simplify this: \[b^m*b^n = b^{m+n}\] \[\frac{b^m}{b^n} = b^{m-n}\] \[b^{-m} = \frac{1}{b^m}\]or\[b^m=\frac{1}{b^{-m}}\] Don't forget that \(b^1 = b\)